OpenStudy (anonymous):

What is the exact volume of a cylinder that has a height of 4.9 yd. and a radius of 2.2 yd. ?

OpenStudy (jdoe0001):

\(\large {\textit{volume of a cylinder}=\pi r^2 h\quad \begin{cases} r=radius\\ h=height \end{cases} }\) plug and chug